OpenStudy (keana):

Find the x value for the point that splits segment CD in half if point C is located at (−2, 4) and point D is located at (3, 7). a. 2.5 b. 1 c. 0.5 d. 3

OpenStudy (anonymous):

add x coordinate of both points and divide by 2 you will get 0.5.

OpenStudy (anonymous):

as you need half the distance so \[\frac{ x_{1} + x_{2} }{ 2 }\] \[=\frac{ -2 + 3 }{ 2 }\] \[=\frac{ 1 }{ 2 } =0\]

OpenStudy (anonymous):

*0.5